In December 1982, a report in the MMWR described three persons who had developed acquired immunodeficiency syndrome (AIDS) but who had neither of the previously known risk factors for the disease: homosexual/bisexual activity with numerous partners and intravenous drug use. These three persons had previously received whole-blood transfusions. By 1983, widespread recognition of the problem of transfusion-related AIDS led to controversial recommendations that persons in known high-risk groups voluntarily defer from donating blood. In June 1984, after the discovery of the human immunodeficiency virus (HIV), five companies were licensed to produce enzyme-linked immunosorbent assay (EIA, then called ELISA) test kits for detecting HIV antibody. A Food and Drug Administration
(FDA) spokesperson stated that, "...getting this test out to the blood banks is our No. 1 priority...." Blood bank directors were anxiously waiting to start screening blood with the new test until March 2, 1985, the date the first test kit was approved by the FDA.

In order to evaluate whether the EIA test should be used as a screening tool, people were tested using the EIA test and had their blood tested to determine whether the HIV antibody was present or not (the gold standard).


   HIV antibody present   HIV antibody absent   Total     
EIA test positive   380   19,992         
EIA test negative   20   979,608         
Total  

Questions   

1. Fill out the rest of the table.

2. Calculate the sensitivity of the EIA test.

3. Calculate the specificity of the EIA test.

4.Calculate the predictive value (+) of the EIA test.
Answer #1

Truth Disease Non Disease (number) (number) Total (number) ТВ Positive (number) Test Positive (True Positive) (False Positive

HIV antibody present HIV antibody absent Total
EIA test positive 380 19992 20372
EIA test negative 20 979608 979628
Total 400 999600 1000000

A=380, B=19992, C=20, D=979608

Sensitivity is the probability that a test will indicate 'disease' among those with the disease:

  • Sensitivity of the EIA Test: A/(A+C) × 100 = 380 / (380 + 20) x 100 = 95%

Specificity is the fraction of those without disease who will have a negative test result:

  • Specificity of the EIA Test: D/(D+B) × 100 = 979608 / (979608+19992) x 100 = 98%

The sensitivity and specificity are characteristics of this test.

The important fact is among the people who test positive, only 1.865% actually have the disease.

  • Positive Predictive Value:A/(A + B) × 100 = 380 / (380+19992) x 100 = 1.865%